Hunter Corp. is considering acquiring Prey Inc. Neither corporation has any debt. Hunter Corp. expects the acquisition of Prey Inc. will generate synergy equal to $1 million per year in after-tax cash flow, indefinitely (a perpetuity, see pages 185 and 186 of the textbook if you need a refresher). Hunter Corp.'s current market value is $70 million. Prey Inc.'s current market value is $30 million. The appropriate discount rate, based on the risk of Prey Inc., is 10%. Hunter Corp. needs to decide between offering 35% of its stock or $37 million in cash to Prey Inc.'s shareholders.

a. What is the value of the target to the acquirer?

b. What is the cost of the stock offer?

c. What is the cost of the cash offer?

d. What are the NPVs of the stock offer and of the cash offer to the shareholders of Hunter?

e. Which alternative should Hunter Corp. pursue? Why?
Answer #1

(a) Market Value of Hunter Corp = A = $ 70 million and Market Value of Pery Inc = B = $ 30 million

Synergy: $ 1million worth of after-tax cash flows every year in perpetuity, Appropriate Discount Rate = 10 %

Therefore, PV (Present Value) of Perpetual Synergy Cash Flows = 1 / 0.1 = $ 10 million

Combined Firm Value = A + B + PV of Synergy = 70 + 30 + 10 = $ 110 million

Gains of the Merger = 110 - Value of Pery Inc - Value of Hunter Corp = 110 - 70 - 30 = $ 10 mllion

(b) Stock Offer = 35 % of Hunter's Stock = 0.35 x 70 = $ 24.5 million

Cost of Merger = Value of Stock Paid - Value of Perry Inc. = 24.5 - 30 = - $ 5.5 million

(c) Cash Offer = $ 37 million

Cost of Merger = Cash Paid - Value of Perry Inc = 37 - 30 = $ 7 million

(d) Merger NPV in Stock Offer:

NPV = Merger Gain - Cost of Merger = 10 - (-5.5) = $ 15.5 million

Merger NPV in Cash Offer:

NPV = Merger Gain - Cost of Merger = 10 - 7 = $ 3 million

(e) Hunter Corp should select the stock offer over the cash offer, as the former has a greater merger NPV for the acquiring firm as compared to the latter.